the difference between two numbers is 28. the larger number is 8 less than twice the smaller number. find both numbers
Answer by JulietG(1812)

A - B = 28
A = 2B - 8
Substitute the value of A
(2B-8) - B = 28
Add 8 to each side
2B - B = 36
B = 36
.
If B is 36, then A is 36+28, or 64
Let's check that.
"the larger number is 8 less than twice the smaller number"
64 = twice 36 less 8
64 = 72 less 8
64 = 72-8 [Valid]
Success!